# Eigen-decomposition of Covariance matrices : An application to the BAO Linear Point

### v1

@@ I Introduction
Some of the tightest constraints on the cosmological distance - redshift relation come from the BAO feature in the pair correlation function Eisenstein et al. (2005); Anselmi et al. (2018a); Alam et al. (2021); Abbott et al. (2024); Adame et al. (2024a). This has led to interest in the precision with which the pair correlation function (TPCF) can be measured, and how this precision translates into uncertainties on the distance scale estimate. As a result, there is significant interest in understanding the covariance between pair counts on different scales.
On the $\sim 140$ Mpc scales of most interest to BAO cosmology, the Gauss-Poisson approximation to the covariance is rather accurate Smith et al. (2008a); Grieb et al. (2016). In this approximation, three different terms contribute: one is a purely cosmological term, the other is a pure shot-noise term, and the third is a combination of the two. The first part of our paper is devoted to a study of the relative importance of these terms. We address this by rotating the covariance matrix into diagonal form, checking how many eigenvectors contribute significantly to the total covariance, and then looking at those eigenvectors. It turns out that this provides a simple way of seeing which term dominates and when, as well as for understanding the shapes of the eigenvectors. We use this insight to explore how binning of the pair counts (width of the rectangular bins, or different bin shapes) affects the structure of the covariance matrix.
@@ II.1 Gauss-Poisson approximation to covariance matrix
Figure 1: Fractional contribution of the sorted eigenvalues $\lambda_{j}$ to the total variance, $\sum_{1}^{i}{\lambda_{j}}/\sum_{1}^{30}{\lambda_{j}}$, for a range of choices of the ratio of ‘signal’ ($P(k)$ amplitude) to ‘noise’ ($1/\bar{n}$). Fiducial and idealized no shot-noise cases are very similar; more than 95% of the total variance comes from the first 3 modes. More modes matter when the noise dominates. Smooth dashed curve shows our analytic expression for the pure noise case: $(2/\bar{n}V)/(\bar{n}V_{i})$. Dotted curve is for $b\to b-1$ but fiducial shot-noise.
@@ II.2 Eigen-decomposition of the covariance matrix
To illustrate our results, we now use Eq. 3 to generate $C_{ij}$, for 30 non-overlapping bins of $\Delta r=2~{}h^{-1}$ Mpc, running from $[60-120]~{}h^{-1}$pc, with the fiducial values of background cosmology, redshift, survey volume, biased tracer number density, and clustering strength mentioned at the start of this section.
@@
The lowermost curve is for the pure shot-noise limit (we have set $P(k)=0$). This case is analytic: $C_{ii}$ is diagonal, with entries $(2/\bar{n}V)/(\bar{n}V_{i})$, where $V_{i}$ is the volume of the $i$th bin. For bins of width $\Delta r=r_{i+1}-r_{i}$ that are equally spaced with spacing $\epsilon\Delta r$ (typically $\epsilon=1$ but we will see later why the more general case is interesting) where $r_{i}$ is the lower bound of the $i$th bin, $V_{i}=(4\pi/3)(\Delta r)^{3}[(j\epsilon-(\epsilon-1)/2)^{3}-(j\epsilon-1-(\epsilon-1)/2)^{3}]$ where $j=i+60/(\epsilon\Delta r)$ and the bin centers are given by $\frac{(r_{i}+r_{i+1})}{2}=60+(i-1)\epsilon\Delta r+\Delta r/2$ since our bins start at $60~{}h^{-1}$pc. Note that all eigenvectors matter. These eigenvectors are delta functions, one for each bin, centered on the middle of the bin.
@@ II.3 Eigen-decomposition of correlation function realizations
where $\xi(s)$ is given by Eq. 2, and the $g_{i}$’s are independent Gaussian random variates with variance $\lambda_{i}$ and mean zero, so the other terms represent the (correlated) scatter around the mean.
@@ II.5 Model-independent error estimates on BAO distance scale
$\bigg{\langle}\frac{\Delta_{\rm{max}}^{2}}{r_{\rm{max}}^{2}}\bigg{\rangle}\approx\sum_{i}\lambda_{i}\,\bigg{(}\frac{r_{\rm max}\Lambda_{i}^{\prime}(r_{\rm{max}})}{r_{\rm{max}}^{2}\xi^{\prime\prime}(r_{\rm{max}})}\bigg{)}^{2}.$ (9)
@@ II.6 Comparison with standard method
We conclude that our methodology is an efficient way of determining accurate uncertainties on the Linear Point estimate of the BAO distance scale. In particular, since our analysis suggests that, for reasonable/realistic values of the shot-noise, the relevant eigenvalues and eigenvectors are entirely determined by the ratio of $P^{2}(k)$ and the survey volume $V_{s}$, they scale as $(b\sigma_{8})^{2}/\sqrt{V_{s}}$. Our curves assumed $(b,\sigma_{8},V_{s})=(1.97,0.84,5~{}h^{-3}{\rm Gpc}^{3})$, so for other values, the fractional error is given by scaling the numbers in our Fig. 5 by $(b_{\rm eff}/2.23)^{2}(\sigma_{8}/0.84)^{2}\sqrt{5/V_{s}}$. This is not quite right, because it ignores the fact that the smearing also depends on $\sigma_{8}$, but it is a useful first guess.
@@ IV Discussion and Conclusions
We argued that, as a result, the cosmic variance eigen-modes alone should provide a good description of the correlated residuals which result from fitting smooth functional forms to the measured TPCF. We provided a simple (Eq. 12) but accurate (Fig. 5) approximation for the uncertainty on the Linear Point estimate of the BAO distance scale which explains why the uncertainty is greater in the evolved field than in linear theory; allows one to quantify the gains from working with the reconstructed signal; and does not depend on the functional form fitted to the TPCF or the binning used. It also provides insight into why the LP is more robust than the inflection point, and why both are more precise distance indicators than the peak or dip scales. Perhaps most importantly, our approximation allows one to quantify the precision of the BAO distance scale estimate without having to generate a large ensemble of mock catalogs. Therefore, it should be useful for estimating the gains in precision which come from making measurements in the reconstructed field (which are often quoted), after marginalizing over the unknown cosmological model (which is often ignored).
The plateaus that are apparent in Figs. 5 and 6 suggest that it should be possible to write down a prescription for determining the optimal number of eigen-modes which should be used in cosmological analyses. In future work, we hope to combine our eigen-mode analysis with the Bayesian framework of Paranjape and Sheth (2022).

### v2

@@ I Introduction
Some of the tightest constraints on the cosmological distance - redshift relation come from the BAO feature in the pair correlation function Eisenstein et al. (2005); Anselmi et al. (2018a); Alam et al. (2021); Abbott et al. (2024a); Adame et al. (2024a). This has led to interest in the precision with which the pair correlation function (TPCF) can be measured, and how this precision translates into uncertainties on the distance scale estimate. As a result, there is significant interest in understanding the covariance between pair counts on different scales.
On the $\sim 140$ Mpc scales of most interest to BAO cosmology, the Gauss-Poisson approximation to the covariance is rather accurate Smith et al. (2008a); Grieb et al. (2016). In this approximation, three different terms contribute: one is a purely cosmological term, the other is a pure shot-noise term, and the third is a combination of the two. The first part of our paper is devoted to a study of the relative importance of these terms. We address this by rotating the covariance matrix into diagonal form, checking how many eigenvectors contribute significantly to the total covariance, and then looking at those eigenvectors. This provides a simple way of seeing which term dominates and when, as well as for understanding the shapes of the eigenvectors. We use this insight to explore how binning of the pair counts (width of the rectangular bins, or different bin shapes) affects the structure of the covariance matrix. Our methodology is similar in spirit to the Karhunen-Loève decomposition Karhunen (1947), in which a stochastic process is represented by orthonormal basis functions and uncorrelated random coefficients. The Karhunen-Loève decomposition was applied to Cosmic Microwave Background (CMB) maps Bond (1995); Bunn and Sugiyama (1995); Bunn (1996) as well as galaxy redshift surveys Vogeley and Szalay (1996) in the 1990s for optimal data compression around the time of the first large-scale cosmological surveys. Tegmark et al. (1997) provide a generalization of this method in anticipation of subsequent generation surveys where the amount of data collected is too large for an uncompressed maximum likelihood analysis. They found that for both CMB maps and redshift-space distortions, a compression by a factor of $\sim 10$ is achievable by keeping the first $\sim 10\%$ of eigen-modes from the covariance matrix.
@@
Although we focus on the Linear Point scale, the scale $r_{0}$ where the pair correlation function crosses zero, $\xi(r_{0})=0$, has also been proposed for use as a standard ruler Prada et al. (2011). Appendix A shows the results of applying our analysis to $r_{0}$.
@@ II.1 Gauss-Poisson approximation to covariance matrix
Figure 1: Fractional contribution of the sorted eigenvalues $\lambda_{j}$ to the total variance, $\sum_{1}^{i}{\lambda_{j}}/\sum_{1}^{30}{\lambda_{j}}$, for a range of choices of the ratio of ‘signal’ ($P(k)$ amplitude) to ‘noise’ ($1/\bar{n}$). Fiducial and idealized no shot-noise cases are very similar; more than 95% of the total variance comes from the first 3 modes. More modes matter when the noise dominates. The smooth dashed curve shows our analytic expression for the pure noise case: $(2/\bar{n}V)/(\bar{n}V_{i})$. Dotted curve is for $b\to b-1$ but fiducial shot-noise.
@@ II.2 Eigen-decomposition of the covariance matrix
To illustrate our results, we now use Eq. 3 to generate $C_{ij}$, for 30 non-overlapping bins of $\Delta r=2~{}h^{-1}$pc, running from $[60-120]~{}h^{-1}$pc, with the fiducial values of background cosmology, redshift, survey volume, biased tracer number density, and clustering strength mentioned at the start of this section.
@@
The lowermost curve is for the pure shot-noise limit (we have set $P(k)=0$). This case is analytic: $C_{ii}$ is diagonal, with entries $(2/\bar{n}V)/(\bar{n}V_{i})$, where $V_{i}$ is the volume of the $i$th bin. For bins of width $\Delta r=r_{i+1}-r_{i}$ that are equally spaced with spacing $\epsilon\Delta r$ (typically $\epsilon=1$ but we will see later why the more general case is interesting) where $r_{i}$ is the lower bound of the $i$th bin, $V_{i}=(4\pi/3)(\Delta r)^{3}[(j\epsilon-(\epsilon-1)/2)^{3}-(j\epsilon-1-(\epsilon-1)/2)^{3}]$ where $j=i+60/(\epsilon\Delta r)$ and the bin centers are given by $\frac{(r_{i}+r_{i+1})}{2}=60+(i-1)\epsilon\Delta r+\epsilon\Delta r/2$ since our bins start at $60~{}h^{-1}$pc. Note that all eigenvectors matter. These eigenvectors are delta functions, one for each bin, centered on the middle of the bin.
@@ II.3 Eigen-decomposition of correlation function realizations
where $\xi(s)$ is given by Eq. 2, the $g_{i}$’s are independent Gaussian random variates with variance $\lambda_{i}$ and mean zero, while $\Lambda_{i}(s)$ are the eigenvectors of $C_{ij}$. Hence, the terms other than $\xi(s)$ represent the (correlated) scatter around the mean.
@@ II.5 Model-independent error estimates on BAO distance scale
$\bigg{\langle}\frac{\Delta_{\rm{max}}^{2}}{r_{\rm{max}}^{2}}\bigg{\rangle}\approx\sum_{i}\lambda_{i}\,\bigg{(}\frac{\Lambda_{i}^{\prime}(r_{\rm{max}})}{r_{\rm{max}}\xi^{\prime\prime}(r_{\rm{max}})}\bigg{)}^{2}.$ (9)
@@ II.6 Comparison with standard method
We conclude that our methodology is an efficient way of determining accurate uncertainties on the Linear Point estimate of the BAO distance scale. In particular, since our analysis suggests that, for reasonable/realistic values of the shot-noise, the relevant eigenvalues and eigenvectors are entirely determined by the ratio of $P^{2}(k)$ and the survey volume $V_{s}$, they scale as $(b\sigma_{8})^{2}/\sqrt{V_{s}}$. Our curves assumed $(b,\sigma_{8},V_{s})=(1.97,0.84,5~{}h^{-3}{\rm Gpc}^{3})$, so for other values, the fractional error is given by scaling the numbers in our Fig. 5 by $(b_{\rm eff}/2.23)^{2}(\sigma_{8}/0.84)^{2}\sqrt{5/V_{s}}$. This is not quite right, because it ignores the fact that the smearing also depends on $\sigma_{8}$, but it is a useful first guess. Additionally, we note that the number of eigen-modes that captures a sufficient amount of information depends on $bP(k)$, $V_{s}$, as well as $\bar{n}$. There are two features in our analysis that show this number: (1) the change of the slope in Fig. 7 where we plot the eigenvalues by mode and (2) the plateaus in Figs. 5, 6, and 9 that show $\sigma_{\rm BAO}$ (which is what we really care about). Although we do not provide an explicit expression here, both features suggest that the first 5 eigen-modes (which contributes to $\sim 95\%$ of the total covariance) are sufficient for our choice of $(b,V_{s},\bar{n})$, which are reasonable values for current and next-generation surveys. While our fiducial choice of binning is 30 non-overlapping bins of $2~{}h^{-1}$pc, our results are rather insensitive to binning - with 5 eigen-modes being sufficient for 15 bins of $4~{}h^{-1}$pc, as well as 10 bins of $6~{}h^{-1}$pc, the reason for which we will see in Section III.
@@ IV Discussion and Conclusions
We argued that, as a result, the cosmic variance eigen-modes alone should provide a good description of the correlated residuals which result from fitting smooth functional forms to the measured TPCF. We provided a simple (Eq. 12) but accurate (Fig. 5) approximation for the uncertainty on the Linear Point estimate of the BAO distance scale which explains why the uncertainty is greater in the evolved field than in linear theory (Fig. 6); allows one to quantify the gains from working with the reconstructed signal; and does not depend on the functional form fitted to the TPCF or the binning used. It also provides insight into why the LP is more robust than the inflection point, and why both are more precise distance indicators than the peak or dip scales. Perhaps most importantly, our approximation allows one to quantify the precision of the BAO distance scale estimate without having to generate a large ensemble of mock catalogs. Therefore, it should be useful for estimating the gains in precision which come from making measurements in the reconstructed field (which are often quoted), after marginalizing over the unknown cosmological model (a step which is often ignored).
Our approach also provides a realistic estimate of the precision of the zero-crossing scale (Fig. 9), showing that the uncertainty on it depends on how steeply $\xi$ crosses zero. Although it is considerably less precise than either $r_{\rm LP}$ or $r_{\rm infl}$ for realistic galaxy surveys, our analysis vastly simplifies the process of quantifying the synergy between joint analyses of $r_{\rm LP}$ and $r_{0}$ in the same dataset.
In practice, our analysis exploits the appearance of extended plateaus in plots of how the precision in the distance scale indicator varies as more (sorted) eigen-modes are included (c.f. Figs. 56 and 9), since it is these plateau heights that correspond to the uncertainty estimates that are returned from fitting basis functions to the measurements. (If a plateau is not seen, then this is an indication that the contributions from the cosmic variance and shot-noise terms to the full uncertainty are not so easily separated, and this makes it harder to pinpoint the correspondence between our analysis and the standard one.) These plateaus suggest that it should be possible to write down a prescription for determining the optimal number of eigen-modes which should be used in cosmological analyses.
While we do not perform cosmological analysis to infer cosmological parameters or evaluate cosmological models in this work, our approach can be used to rapidly measure the BAO scale and its uncertainties accurately given a dataset. This will enhance the utility of the LP that previous works have already shown: (1) the distance scale obtained from the LP is consistent with the standard template-fitting method, while being more model agnostic Anselmi et al. (2018a) and (2) the resulting cosmological parameters inferred from the LP are consistent with the standard approach with 20 to 30% larger uncertainties under a flat $\Lambda$CDM cosmology He et al. (2023). Furthermore, the rapid estimation of the BAO scale and its uncertainties means that performing multiple BAO analyses assuming a variety of different cosmological models will be possible at reasonable computational cost, which will be especially important in the light of the time-evolving dark energy results from Abbott et al. (2024b); Adame et al. (2024a). In future work, we hope to combine our eigen-mode analysis with the Bayesian framework of Paranjape and Sheth (2022) utilizing the Laguerre reconstruction methodology of Nikakhtar et al. (2021a, b), for a completely model-agnostic BAO analysis.
@@ Acknowledgements.
Appendix A Application to the zero-crossing scale
With some care, the zero-crossing scale, $r_{0}$, can also be used as a standard ruler Prada et al. (2011); Gaines et al. (2024). In practice, estimating $r_{0}$ is similar to the Linear Point scale: one fits the pair counts with suitably chosen basis functions before estimating the zero-crossing from the fit. Therefore, the analysis in the main text can be modified to provide estimates of the uncertainty on $r_{0}$ as follows.
If $s_{0}$ is the scale on which $\Xi(s)=0$, and $r_{0}$ is where $\xi(r)=0$, we have
$0=\xi(r_{0})+\Delta_{0}\,d\xi/dr+\sum_{i}g_{i}[\Lambda_{i}(r_{0})+\Delta_{0}\,d\Lambda/dr]$ (21)
where $\Delta_{0}=s_{0}-r_{0}$. Since $\xi(r_{0})=0$, dividing throughout by $\xi^{\prime}(r_{0})$ and then assuming $\Lambda^{\prime}(r_{0})/\xi^{\prime}(r_{0})\ll 1$, yields
$\frac{\Delta_{0}}{r_{0}}\approx-\sum_{i}g_{i}\,\frac{\Lambda_{i}(r_{0})}{r_{0}\xi^{\prime}(r_{0})},$ (22)
so
$\left\langle\frac{\Delta_{0}^{2}}{r_{0}^{2}}\right\rangle\approx\sum_{i}\lambda_{i}\,\left[\frac{\Lambda_{i}(r_{0})}{r_{0}\xi^{\prime}(r_{0})}\right]^{2}.$ (23)
Notice that the uncertainty is smaller if $\xi^{\prime}$ is bigger. This is sensible: the zero-crossing of a curve is easier to detect if the curve is steep. In practice, this means that more biased tracers are to be preferred for this measurement.
Figure 9: Same as Figure 5, but now showing the fractional uncertainties on the zero-crossing scale (grey) as well.
Figure 9 shows how well this compares to the standard approach (of fitting a polynomial to the pair counts over the range $100-160~{}h^{-1}$pc in bins of width $\Delta r=2~{}h^{-1}$pc) as the number of modes that is included in the sum increases. The figure is in the same format as Figure 5 in the main text, to highlight the similarity of the analysis. The height of the plateau region defined by the grey symbols is our eigen-mode estimate of the uncertainty on $r_{0}$. It tracks that from the standard (polynomial fitting) method quite well (grey dashed), showing that it provides reliable estimates for analyses of the zero-crossing scale.
For $r_{0}$, there are two striking differences compared to our analyses of the scales ($r_{\rm LP}$ etc.) that we highlighted in the main text. First, the plateau region – which the main text argued is a proxy for the uncertainty on the distance scale – is reached after only the first couple of eigen-modes. For $r_{\rm LP}$ etc., more eigen-modes are need before convergence to a plateau is seen, and then the plateau lasts for smaller range of eigen-modes before the variance increases again (mainly due to shot-noise dominated modes). Second, the uncertainty on $r_{0}$ is much larger than on $r_{\rm LP}$ etc. This quantifies results which are obvious by eye in Ref. Gaines et al. (2024). In this respect, $r_{\rm LP}$ is the better distance scale indicator than $r_{0}$. Nevertheless, there are synergies associated with having two distance scale estimates from the same measurement which we will explore elsewhere.